OpenStudy (anonymous):

Of 50 employees of a store located in downtown Boston, 18 people take the subway to work, 12 take the bus, and 7 take both the subway and the bus. How may employees (a) take the subway or the bus to work? (b) take only the bus to work? (c) take either the bus or the subway to work? (d) get to work by some other means?

OpenStudy (anonymous):

a-47 b-5 c-16 d-10

OpenStudy (anonymous):

sorry a is 40

OpenStudy (anonymous):

can you tell me how you got those numbers please

OpenStudy (anonymous):

For a, I id 50-12+7=45 for b, 50-18+7 = 39 For c, 50-7=43 For d, 50-18-12-7=13

OpenStudy (anonymous):

you have to make a venn diagram for solving it I am uploading the diagram, pls wait

OpenStudy (anonymous):

Oh great, thanks I thought i lost you on here..lol

OpenStudy (anonymous):

here it is lol inside the circles there are the no. of people who only take the bus or subway and in the intersection of the circle are the people who take both of them.

OpenStudy (anonymous):

did you understand it yet?

OpenStudy (anonymous):

Yes, I do; thank you so very much

OpenStudy (anonymous):

Thanks to the diagram...it really helps..

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

Okay, it was very helpful, but for some reason I can't see how you got the other answers..i really want to understand..brain not working

OpenStudy (anonymous):

umm, okay in the question it tells you the no. of people going from bus and subways right like for bus-12 and subway-18 right? it is telling you how many people used bus and subway but it isn't telling how many of them used both

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

the no. of people who used both is given 7 so you need to deduct 7 from the no. of people who traveled from bus and subway that is, 18-7=11 and 12-7=5

OpenStudy (anonymous):

because the people who went with both the transport ways got mixed up with the people who went from only one of them so you need to subtract them from the nos. gives to you

OpenStudy (anonymous):

just see the diagram, the people who went only from bus are 5 and the no. of people who only went from subway are 11

OpenStudy (anonymous):

please do tell me if you didn't understand

OpenStudy (anonymous):

Okay, i'm trying here

OpenStudy (anonymous):

According to the diagram, (b) take only the bus would be 11

OpenStudy (anonymous):

Take only subway would be 5

OpenStudy (anonymous):

7 would be (c) either

OpenStudy (anonymous):

And (d) 50-18-12=20

OpenStudy (anonymous):

(c) would be 11+5=16

OpenStudy (anonymous):

and you miscalculated (d)

OpenStudy (anonymous):

13? 50-18-12-7=13

OpenStudy (anonymous):

nono you did it right but 50 -12-18 is 10

OpenStudy (anonymous):

My calculator says 20 which is why i'm confused

OpenStudy (anonymous):

oh yeah i m really sorry my bad :P

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

12+18=30, then subtract from 50 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Lol..wow, i actually got it..lol I really apreciate you helping me so much..

OpenStudy (anonymous):

i m really glad that you got it

OpenStudy (anonymous):

Can I ask you another question since your so good at this I just need to verify something

OpenStudy (anonymous):

If i post it on another one to get you another metal

OpenStudy (anonymous):

lol theres no need for that

OpenStudy (anonymous):

In a survey conducted by a union, members were asked to rate the importance of the following issues: (1) job security, (2) increased fringe benefits, and (3) improved working conditions. Five different responses were allowed for each issue. Among completed surveys, how many different responses to this survey were possible?

OpenStudy (anonymous):

Thank you...The answer I got was 5*3=15 different responses..assuming 5 different responses for all 3